OpenStudy (hannaha):

How do you find the square root of 121? Medal for fast help! :)

OpenStudy (anonymous):

Answer: 11

OpenStudy (hannaha):

thanks, any way you could summarize how to figure that out?

OpenStudy (anonymous):

first find all odd or even prime factors for given no. then on applying square root we get the required value

OpenStudy (hannaha):

what does applying square root mean?

OpenStudy (hannaha):

1, 11 and 121 are the prime factors of 121

OpenStudy (anonymous):

Just now that 10 * 11 is 110 so if you add 11 more you get 121 so therefore you now that the answer is 11 * 11

OpenStudy (hannaha):

okay so basically square root is what number times itself can equal the given number?

OpenStudy (zarkon):

the square root of 121 is 11 and -11 \(b\) is the square root of a number \(a\) if \(b^2=a\) since \(11^2=121\) and \((-11)^2=121\) both 11 and -11 are square roots of 121

OpenStudy (hannaha):

Okay :) Thanks!